### v1

Abstract
Context: Deep learning has achieved remarkable progress in various domains. However, like any software system, deep learning systems contain bugs, some of which can have severe impacts, as evidenced by crashes involving autonomous vehicles. Despite substantial advancements in deep learning techniques, little research has focused on reproducing deep learning bugs, which hinders resolving them. Existing literature suggests that only 3% of deep learning bugs are reproducible, underscoring the need for further research.
Objective: This paper examines the reproducibility of deep learning bugs. We identify edit actions and useful information that could improve deep learning bug reproducibility.
Method: First, we construct a dataset of 668 deep learning bugs from Stack Overflow and Defects4ML across three frameworks and 22 architectures. Second, out of the 668 bugs, we select 102 bugs using stratified sampling and try to determine their reproducibility. While reproducing these bugs, we identify edit actions and useful information necessary for their reproduction. Third, we used the Apriori algorithm to identify useful information and edit actions required to reproduce specific bug types. Finally, we conduct a user study with 22 developers to assess the effectiveness of our findings in real-life settings.
Results: We successfully reproduced 85 out of 102 bugs attempted. We identified ten edit actions and five useful information categories that can help us reproduce the deep learning bugs. Our findings improved bug reproducibility by 22.92% and reduced reproduction time by 24.35% based on a user study with 22 developers.
Conclusions: Our research addresses the critical issue of deep learning bug reproducibility. Practitioners and researchers can leverage our findings to improve deep learning bug reproducibility.
1 Introduction
Deep learning (hereby DL) has been widely used in many application domains, including natural language processing, finance, cybersecurity med1 ; fin1 ; cyb1 , autonomous vehicles pei2017 ; ma2018 , and healthcare systems esteva2019guide .
Like any software system, deep learning systems are prone to bugs. Bugs in deep learning systems could arise from various sources, including errors in the dataset, incorrect hyperparameters, and incorrect structure of the deep learning model BRAIEK2020110542 . These bugs could lead to program failures, poor performance, or incorrect functionality, as reported by existing literature islamfse19 . They can also lead to serious consequences, as shown by the fatal crash involving Uber’s self-driving car selfdrivingcarcrash . Thus, we must fix the bugs before deploying a deep learning model in production. However, one must reproduce the bugs before fixing them to verify their presence/absence in the system. Unfortunately, reproducing DL bugs is challenging due to deep learning systems’ multifaceted nature and dependencies, which encompass data, hardware, libraries, frameworks, and client programs. Furthermore, deep learning systems are inherently non-deterministic (i.e., random weight initialization), which leads to different outcomes across multiple runs and thus makes the reproduction of deep learning bugs challenging nagarajan2018impact . Moreover, they also suffer from a lack of interpretability krishnan2020against , which makes the reproduction of deep learning bugs challenging in comparison to traditional software bugs. According to existing investigations defects4ml , only 3% of DL bugs are reproducible, further demonstrating the challenges in reproducing DL bugs.
Existing literature investigates the challenges of reproducing programming errors or bugs from various sources. Mondal et al. msr2019mondal investigate the challenges in reproducing programming issues reported on Stack Overflow and suggest several edit actions to help reproduce them. Rahman et al. icsme2020rahman conduct a multi-modal study to understand the factors behind the non-reproducibility of software bugs. They identify 11 significant factors behind bug non-reproducibility, including missing information, bug duplication, false positive bugs, and intermittency. Overall, these studies highlight the challenges in reproducing traditional software bugs but do not deal with any deep learning bugs, which warrants further investigation.
Recently, a few studies have attempted to tackle the challenges of deep learning bugs. Liang et al. liang2022gdefects4dl provide a dataset of 64 deep learning bugs collected from GitHub issues. They classify these bugs into six categories according to the taxonomy of Humbatova et al. taxonomyRealFaults . Moravati et al. defects4ml constructed a benchmark dataset containing 100 deep learning bugs collected from StackOverflow and GitHub; they reproduced each of them. Although these studies offer benchmark datasets containing reproducible bugs, their primary focus is dataset construction. They do not report any detailed instructions (e.g., sequence of actions) essential to the reproduction of the deep learning bugs. Our work attempts to fill this important gap in the literature.
In this paper, we conduct an empirical study to better understand the challenges in reproducing deep learning bugs. First, we collect a total of 568 DL bugs from Stack Overflow posts. Then, we extend them with 100 DL bugs from the benchmark dataset of Moravati et al. defects4ml , which makes up our final dataset of 668 DL bugs. Using the taxonomy of Humbatova et al. defects4ml , we divide these bugs into five categories: 167 model, 213 tensor, 145 training, 113 GPU, and 30 API bugs. Second, by using stratified sampling, we select 102 of these bugs and determine the reproducibility status of these bugs by attempting to reproduce them using the code snippet and complementary information from Stack Overflow and the benchmark dataset. Third, the first author manually analyzed the produced artefacts to determine the categories of information that are useful for deep learning bug reproduction. Fourth, we use the information gathered during the bug reproduction and create a dataset of transactions that link the type of bug with edit actions and information categories. Then, our study establishes connections among the information categories, edit actions, and the type of bugs by employing the Apriori algorithm apriori . Finally, to further validate our findings, we conducted a developer study with 22 participants from industry and academia. Half of the participants were asked to reproduce bugs using our identified information, while the other half were asked to reproduce the same bugs without access to that information. Results from the user study show that our recommended edit actions and information categories can reduce the time to reproduce the bug by 24.35%. Thus, we answer three important research questions as follows:
RQ1: Which edit actions are useful for reproducing deep learning bugs?
Determining the key edit actions that are crucial for reproducing deep learning bugs is important since almost none of the bugs can be reproduced using the verbatim code snippet. By manually reproducing 85 deep learning bugs, we identify ten key edit actions that could be useful to reproduce deep learning bugs (e.g., input data generation, neural network construction, hyperparameter initialization). -
RQ2: What information categories and edit actions are useful for reproducing specific types of deep learning bugs?
Different types of DL bugs need different types of information or edit actions, which warrants further investigation. Using the Apriori algorithm, we have determined the top 3 pieces of information and the top 5 edit actions that can help one reproduce each type of bug. These insights can be used not only to detect the missing information in a submitted bug report but also to formulate the follow-up questions soliciting the missing information. -
RQ3: Do the suggested edit actions and identified information categories contribute to enhancing the reproducibility of deep learning bugs in practice?
To assess the effectiveness of our suggested edit actions and identified information categories in improving bug reproducibility, we conducted a user study involving 22 professional developers. In our developer study, the participants assigned to the experimental group used our suggested edit actions and information categories to reproduce deep learning bugs. In contrast, the participants in the control group reproduced the bugs without access to the suggested edit actions and information categories. Based on the feedback from the developers, we found that our recommended edit actions and highlighted information categories (a) helped them reproduce 22.92% more bugs and (b) decreased their time to reproduce the deep learning bugs by 24.35%.
The remainder of the paper is organized as follows. Section 2 presents a motivating example that highlights the challenges in reproducing deep learning bugs using existing techniques. Section 3 then describes our study methodology in detail, including data collection, dataset construction, environment setup, qualitative analysis, and user study design. Next, Section 4 presents the findings by answering our three research questions. It summarizes the key edit actions, useful information categories, and their relationships with bug types that enhance deep learning bug reproducibility. Section 5 and Section 6 provide a discussion of our results, including a proof-of-concept tool, comparison with language models and threats to validity. Finally, Section 7 reviews related literature, and Section 8 concludes the paper.
2 Motivating Example
Bug reports or programming Q&A posts might not always provide sufficient information to reproduce a deep learning bug. Let us consider the example question shown in Fig. LABEL:fig:irreproducibleMotivatingExample motivatingExample1 . Here, the user attempts to train a neural network using Keras. The input data type is a object, and the user aims to pass the training dataset to the method. However, s/he discovers that their object is not recognized by the Keras library. The user also provides a code snippet to aid the reproducibility of the bug. Unfortunately, the issue cannot be reproduced since the provided information does not contain any dataset for training, network definition, hyperparameters, or required dependencies. In Stack Overflow, this question has failed to receive a precise response. Even though the above code (Fig. LABEL:fig:irreproducibleMotivatingExample) could be made compilable or executable using edit actions, the bug cannot be reproduced due to its complex nature. Since the earlier study msr2019mondal does not deal with any deep learning bugs, their suggested edits might also not be effective.
Let us consider another example question shown in Fig. LABEL:fig:motivatingExample motivatingExample2 . Here, the user wants to obtain the confusion matrix from a multi-class classification model. Unfortunately, s/he runs into a runtime error, as the method does not support multiple output labels. To reproduce the bug, we first determine the type of this bug. By manually analyzing the Stack Overflow post, code snippet, and the corresponding stack trace, we found that the error was related to the incorrect shape of data being passed to the method. Since the bug was triggered by a wrong shape of the training dataset, it falls into the ‘Tensor’ category according to the taxonomy of Humbatova et al. taxonomyRealFaults . Then, we check the useful information suggested by our study (see RQ2) for reproducing tensor bugs. The top 3 items are Data, Logs, and Code Snippet, as highlighted in Table. 7. Although the question in Fig. LABEL:fig:motivatingExample contains the relevant code snippet and logs, it does not contain any information about the dataset for training, which is the most useful information for reproducing tensor bugs according to our findings (RQ2). Subsequently, we also check the recommended edit actions to reproduce a tensor bug. From Table 8, we use the top 3 edit actions – Input Data Generation, Import Addition, and Hyperparameter Initialization. When these edit actions and information categories were applied to the code, we were able to successfully reproduce the bug, which indicates the effectiveness of our suggested actions and information. Similarly, this issue was reproduced by other users of Stack Overflow, and the question received a correct solution within two hours of its submission, as shown in Fig. LABEL:fig:AcceptedAnswer.
3 Study Methodology
Fig. 1 shows the schematic diagram of our empirical study. We discuss different steps of our study as follows.
3.1 Selection of Data Sources
We select Stack Overflow as a primary data source for our study. It is the largest programming Q&A site for programming topics containing over 24 million questions and 35 million answers SOUsageData . Thus, Stack Overflow could be a potential source for deep learning (DL) bugs. Developers often submit their encountered problems on Stack Overflow, when building their deep learning applications. According to a recent work zhao2021state , Stack Overflow contains at least 30 topics and 80K posts related to deep learning issues, which makes it a potential source for our data. We also select the benchmark dataset of Moravati et al. defects4ml containing 100 DL bugs as our second source of data.
To gather relevant posts from Stack Overflow, we use the Stack Exchange Data Explorer platform111https://data.stackexchange.com/stackoverflow/query/new. We employ multiple filters to capture relevant posts discussing deep learning bugs. First, we select the recent posts (i.e., submitted between May 2020 and May 2023) with the following tags: ‘tensorflow’, ‘keras’, and ‘pytorch’. These tags were selected since they represent the most frequently used frameworks for deep learning zhang2019empirical . This filtration resulted in 14,065 posts for PyTorch, 14,971 posts for Tensorflow, and 3,152 posts for Keras. We then applied the following four filtration criteria to remove conceptual questions from our collection:
Keyword Filtering: This filtration removes the how-to questions and the questions requesting installation instructions. To filter these questions, we use appropriate keywords (‘how’, ‘install’, and ‘build’), as recommended by Humbatova et al. taxonomyRealFaults .
Code Snippet: Given our focus on the reproducibility of reported issues, the questions must include code segments. Therefore, we consider such questions that contain at least one line of code, as suggested by Mondal et al. msr2019mondal .
Accepted Answer: We only select the questions with accepted answers, ensuring that each reported issue (e.g., DL bug) was reproduced and fixed, as advised by Humbatova et al. taxonomyRealFaults .
Negative Question Score: This filtration helps us discard questions that the community has found to be of low quality, as proposed by Ponzanelli et al. ponzanelli2014 .
We combine all four filtration criteria above, construct a SQL query, and then execute the query against the Stack Exchange Data Explorer. The SQL query is available in our replication package replicationPackage . After this filtration step, we obtained a total of 279 posts for Keras, 1,433 posts for Tensorflow, and 1,700 posts for PyTorch. To ensure the validity of our chosen filtration criteria and that the posts are related to DL bugs, we conducted a follow-up manual analysis on a representative sample of 75 bugs. We select the sample for the manual analysis using a 90% confidence interval and a 10% margin of error. We analyze the entire Stack Overflow discussion and code changes in GitHub for each of these bugs, and we found that 96% of them discuss deep learning bugs, confirming the validity of our filtration criteria (Steps 1-2, Fig. 1).
3.2 Dataset Construction
The above filtration step resulted in 3,412 Stack Overflow posts. To construct our final dataset, we apply another filter to these posts where we leverage the tags from Humbatova et al. taxonomyRealFaults . Our idea was to select representative samples from different types of deep learning bugs. To reduce the risk of miscategorization, we thus only selected the tags in the taxonomies and sub-taxonomies proposed by Humbatova et al. Table. 1 shows our selected 3 tags for model bugs, 1 for tensor bugs, 11 for training bugs, 3 for GPU bugs, and 7 for API bugs. We look for these tags in the above Stack Overflow posts and collect 113 Model bugs, 193 Tensor Bugs, 95 Training Bugs, 101 GPU Bugs, and 24 API Bugs. We also discover 42 bugs belonging to multiple categories (e.g., 5 Model & Tensor Bugs, 5 Tensor & GPU Bugs, 3 Training & API Bugs). Thus, the final dataset contains a total of 668 bugs (568 from our dataset + 100 from Defects4ML defects4ml ) and captures a balanced representation from different types of DL bugs, as shown in Table 2 (Step 3, Fig. 1).
| Type of Bug | Tags |
| Model | layer, model, activation-function |
| Tensor | tensor |
| Training | loss-function, training-data, optimization, loss, |
| data-augmentation, performance, learning-rate, | |
| hyperparameters, initialization, imbalanced-data, nan | |
| GPU | gpu, nvidia, cuda |
| API | typeerror, valueerror, attributeerror, importerror, |
| compilererrors, syntaxerror, modulenotfounderror |
| Type of Bugs | Total Number of Bugs |
| Model | 113 |
| Training | 95 |
| GPU | 101 |
| API | 24 |
| Tensor and Input | 193 |
| Mixed | 42 |
| Total | 568 |
3.3 Environment Setup
For our experiments, we use the following environment setup:
Code Editors: We use Visual Studio Code v1.79.0222https://code.visualstudio.com/ and PyCharm 2023.1.1 333https://www.jetbrains.com/pycharm/ to execute code snippets and reproduce bugs from our dataset. Visual Studio Code and PyCharm are popular code editors for building DL-based applications GeeksforGeeks_ide_dl .
Dependencies: To detect the API libraries adopted by the code snippets, we use the pipreqs package444https://pypi.org/project/pipreqs/. We also install the dependencies for each bug into a separate virtual environment using the venv555https://docs.python.org/3/library/venv.html module.
Frameworks: Since our dataset contains bugs from Tensorflow, Keras, and PyTorch, we used all three frameworks in our experiments.
Libraries: For generating the random inputs and visualizing the training metrics, we leveraged several scientific computing libraries such as numpy, pandas, and matplotlib.
Python Version: During our experimentation, we used Python v3.10 to reproduce the deep learning bugs, the latest stable version. If the Stack Overflow issue reported a Python version other than v3.10, we reproduce the bug using the mentioned version instead.
Hardware Config: Our experiments were run on a desktop computer having a 64-bit Windows 11 Operating System with 16GB primary memory (i.e., RAM) and 8GB GPU Memory (Intel(R) Iris XE Graphics).
3.4 Qualitative Analysis
We follow a two-step approach to reproduce the deep learning issues reported in the studied Stack Overflow posts. First, we gather complementary information about the issue, such as the dataset, library versions, and the framework used. Second, we attempt to reproduce the issue using the code snippet and supporting data (e.g., dataset information, environment configurations, hyperparameters, and training logs). Following the successful reproduction of a bug, we capture various information related to each bug, such as the deep learning architecture involved, edit actions used to reproduce the issue, time taken to reproduce the issue, type of issue reproduced, and the type of information present in the Stack Overflow post. All these data gathered during issue reproduction will help us determine the useful information and edit actions needed to reproduce specific types of deep learning bugs (Step 4, Fig. 1).
| Edit Action | Overview |
A: Input Data Generation (D) |
Generating input data which simulates the data used for training the model. |
A: Neural Network Construction (N) |
Reconstructing or modifying the neural network based on the information provided. |
A: Hyperparameter Initialization (H) |
Initializing the hyperparameters for training, such as batch size and number of epochs. |
A: Import Addition and Dependency Resolution (R) |
Determining the dependencies in the code snippet and adding the missing import statements. |
A: Logging (L) |
Adding appropriate logging statements to capture relevant information during reproduction. |
A: Obsolete Parameter Removal (O) |
Removing outdated parameters or functions to match the parameters of the latest library versions. |
A: Compiler Error Resolution (C) |
Debugging and resolving compiler errors that arise due to syntactic errors in the provided code snippet. |
A: Dataset Procurement (P) |
Acquiring the necessary datasets and using them to train the model. |
A: Downloading Models & Tokenizers (M) |
Fetching pre-trained models and tokenizers from external sources. |
A: Version Migration (V) |
Updating code to adapt to changes introduced in newer library or framework versions. |
3.5 Identifying Type Specific Information and Edit Actions
3.5.1 Algorithm Selection
To establish a relationship among the bug types, information categories, and the key editing actions necessary for their reproduction, we use the Apriori apriori algorithm. Apriori is a well-known algorithm for mining frequent item sets from a list of transactions. It helps one identify common patterns and associations between different elements. In our context, we apply the Apriori algorithm to analyze the information gathered during bug reproduction. Our goal was to determine the frequent combinations of bug types, information categories, and edit actions during bug reproduction. We leveraged the Apriori algorithm’s systematic pattern-mining capabilities to establish these relationships.
3.5.2 Generating the Transactions
To create our datasets for the Apriori algorithm, we employed a character encoding, where we encoded all labels into a unique character (e.g., ‘Training Bug’ was encoded as ‘T’, ‘Model Bug’ was encoded as ‘M’, ‘Obsolete Parameter Removal’ was encoded as ‘O’) and converted the data into transactions using the following format.
Example:
Description: The transaction above indicates that the reproduced bug is a training bug (T). The corresponding bug report contains useful information about the bug, such as the dataset used for training (D) and hyperparameters used by the model (H).
Example:
Description: The transaction above indicates that the reproduced bug is a model bug (M). To reproduce the bug, we performed three edit actions, as described below:
Obsolete Parameter Removal (O): We removed some of the parameters that were absent in the recent library and framework version to ensure that the code compiles.
Logging (L): We logged various intermediate program states to verify the bug’s presence.
Neural Network Definition (N): We reconstructed the neural network based on the information provided in the bug report.
Following the specified format, we created two datasets of transactions that associate bug types with crucial information and edit actions, respectively. After creating the transactions, we use the Apriori algorithm to compute the support and confidence for our generated itemsets and association rules. We talk about these metrics in detail below.
3.5.3 Metrics for Apriori Algorithm
Support is the proportion of transactions in the dataset that contain a particular item set. Mathematically, the support of an itemset X is defined as the ratio of the number of transactions containing X to the total number of transactions. It is expressed as:
For a rule , where and are two itemsets, the support is calculated for the combined itemset .
Confidence measures the likelihood that a rule holds. It is defined as the ratio of the support of the combined itemset to the support of the antecedent itemset . Mathematically, confidence is expressed as:
Confidence values range from 0 to 1. A high confidence value indicates a strong association between antecedents and consequent itemsets.
3.5.4 Association Rule Generation
We conducted two separate association rule mining operations in our study - one focused on information categories while the other focused on edit actions used to reproduce deep learning bugs. In particular, we extracted 27 itemsets and 34 rules for information categories, highlighting the useful information for reproducing deep learning bugs. Similarly, we extracted 126 item sets and 284 rules for edit actions, capturing the edit actions to reproduce bugs. We generated association rules based on the entries in our dataset and did not filter or remove any rules before determining confidence values. We then calculate the confidence values for all generated rules to identify the most influential ones for connecting bug types with edit actions and useful information.
3.5.5 Computation of Confidence Values for the Generated Association Rules
As discussed earlier, confidence indicates how frequently a rule occurs, while support indicates the generality of the rule. To compute the confidence values for each association rule, we performed the following steps:
Calculate Support for Antecedent (): We calculate the support for the antecedent, which is the bug type in our case (e.g., ‘T’ for Training Bug or ‘M’ for Model Bug). Support for is the proportion of all transactions that contain the specific bug type.
Calculate Support for Combined Itemset (): We then calculate the support for the combined itemset, , which includes both the bug type and the information category or edit action (e.g., ‘T H’ for Training Bug associated with Hyperparameter Information, or ‘M O’ for Model Bug associated with Obsolete Parameter Removal).
Compute Confidence for the Rule (): The confidence of the rule is computed by dividing the support of the combined itemset by the support of the antecedent . This step gives us the confidence value, which indicates how often the information category or edit action is associated with the bug type in our transactions.
For example, consider the rule . This rule indicates that if the type of bug is a ‘Training Bug’ (T), it can be reproduced by the edit action ‘Input Data Generation’ (D). To calculate the confidence of this rule, we count the number of transactions in which the training bug is reproduced by using the edit action ‘Input Data Generation’ and divide this count by the total number of transactions involving training bugs in the dataset.
3.5.6 Identification of High Confidence Associations
We use high confidence and support values to detect the rules that reliably capture the core factors necessary to reproduce specific types of deep learning bugs. Based on these high-confidence rules, we identify the top 3 pieces of useful information and the top 5 edit actions used to reproduce each bug type. The decision to select three useful pieces of information and five edit actions was influenced by two key factors. First, we adhere to the Parsimony Principle gori2023machine , which suggests that selecting the simplest set of rules is preferable when multiple rules can predict or describe the same phenomenon. We thus concentrate on the most significant factors by selecting the top 3 and top 5 rules for edit actions and useful information, respectively. Second, we filter the rules based on minimum confidence values of 30%, as suggested by Liu et al. liu1999mining . With our limited dataset, a 30% confidence threshold can indicate a substantial pattern since finding associations in 30% of cases points to a meaningful correlation given the data size. Furthermore, the 30% minimum confidence helps filter out spurious correlations with small datasets that can occur by chance. Therefore, for our dataset, the selected threshold strikes an effective balance - it is high enough to identify meaningful associations in the data while eliminating noise from false correlations. Overall, this filtration left us with 23 rules for edit actions and 20 for useful information. Focusing on these high-confidence, high-support rules can reveal the patterns that reproduce deep learning bugs (Step 5a, 5b, Fig. 1).
3.6 User Study
To assess the benefits and implications of our findings in a real-life setting, we conduct a user study involving 22 users (10 from academia + 12 from industry) (Step 6, Fig. 1). We discuss our study setup, including instrument design and participation selection, as follows:
Instrument Design: We divided our questionnaire into three sections. We discuss them in detail below:
Introduction: We first summarize our findings on the reproducibility of deep learning bugs to provide the developers with the necessary context and background information.
Demographic Information: After providing the contextual information about our study, we collect demographic information from developers (e.g., experience bug fixing in deep learning frameworks). We then ask the developers to elaborate on the challenges that they face when reproducing deep learning bugs in their daily lives.
Questionnaire Preparation: First, we select eight bugs (2 Tensor, 2 API, 2 Model, and 2 Training Bugs) from our dataset. Second, we prepare four sets of 2 bugs each by randomly selecting bugs from those eight bugs. Third, we provide the users with the issue description and the code snippet from the original Stack Overflow post. We also provide a Google Colaboratory notebook containing the code for sample edit operations to aid the bug reproduction process. Finally, we include a free-text box in the form to allow the users to share any additional information about edit actions not covered in our study.
Study Session: During our study session, each participant completes the following five tasks. First, the participant provides their demographic information. Second, the participant explains their daily challenges when reproducing deep learning bugs. Third, the participant reproduces two deep learning bugs and reports the edit actions and information they used to reproduce the bugs. Fourth, the participant also provides the rationale behind their used edit actions and information categories to reproduce the bugs. Finally, the participant provides information about any other edit actions that they might have used to reproduce the bugs but are not covered in our study.
| DL Experience | Developers in Control Group | Developers in Experimental Group |
| <1 Year | 3 (27.27%) | 3 (27.27%) |
| 1-5 Years | 6 (54.54%) | 5 (45.45%) |
| 5-10 Years | 2 (18.18%) | 3 (27.27%) |
Participant Selection: We first conducted a pilot study with two researchers and two developers. Based on their feedback, we rephrased ambiguous questions and added sample code to aid the manual bug reproduction by the users. Incorporating this constructive feedback enabled us to refine and improve the quality of our final questionnaire. Then, we invite professional developers and researchers with relevant deep learning experience to our study. We send our invitations to the potential participants using direct correspondences, organization mailing lists (e.g., Mozilla Firefox), and public forums (e.g., LinkedIn and Twitter). A total of 22 participants responded to our invitations. Out of them, 10 (45%) came from academia, and 12 (55%) came from industry. In terms of bug-fixing experience with deep learning bugs, 14 participants (63.63%) had 1-5 years of experience, 4 (18.18%) had 5-10 years of experience, and the remaining 4 (18.18%) had less than one year of experience. In terms of deep learning frameworks, 19 participants (86.34%) reported having working experience with Tensorflow, 21 (95.45%) reported experience with PyTorch, and 20 (90.91%) reported experience with Keras. All these statistics indicate a high level of cross-framework expertise within our participants. We carefully divided the developers into control and experiment groups, as shown in Table. 4. Our guiding principle was to ensure that both groups had a similar distribution in their relevant experience. Specifically, we surveyed all the developers about their experience and used a stratified random sampling approach to assign developers to the two groups. This randomization allowed us to minimize potential bias and confounding factors across the groups. Using these two groups, we wanted to assess the benefit of our recommended information in the context of bug reproduction. The control group receives no hints about how to reproduce the bug. In contrast, the experimental group receives hints (e.g., useful information, edit actions) that could help them reproduce a bug.
| Type of Bug | Bugs Reproduced | Model Architectures Covered |
| Training (T) | 26 | CNN, LSTM, AutoEncoder, MLP, RCNN, |
| ResNet | ||
| Model (M) | 20 | BERT, CNN, GMM, LSTM, MLP |
| VGG16, Transformers | ||
| API (A) | 11 | CNN, GAN, MLP, Transformers, VGG19, |
| Variational RNN | ||
| GPU (G) | 3 | - |
| Tensor and Input (I) | 21 | CNN, GAN, Logistic Regression, MLP, ResNet |
| Mixed (X) | 4 | CNN, BERT, MLP |
3.7 User Study Results Analysis
We perform an agreement analysis between our empirical findings and the findings from the user study to assess their generalizability and strengthen our understanding of the reproducibility of deep learning bugs. To perform the agreement analysis, we compare the edit actions used in our manual bug reproduction process with those used by participants in the user study. If similar edit actions were used, it would indicate a strong agreement between the empirical findings and user study results.
In our user study, the participants first reproduce their assigned bugs and report their used information or actions to reproduce their bugs. To analyze the effectiveness of our recommended information, we compare the control and experimental groups in terms of their success rates in bug reproduction and the time taken to reproduce the bugs. We select bug reproducibility rate and time taken for bug reproduction as our key metrics to evaluate the effectiveness of our findings. These quantitative metrics directly measure how successful and efficient our recommended information is in assisting the developers to reproduce deep learning bugs. Higher reproducibility rates and reduced reproduction times in the experimental group compared to the control group would indicate that our findings are effective for improving deep learning bug reproducibility. We also analyze the qualitative feedback from the participants to determine if provided hints were useful or not. We also analyze them to uncover new insights into deep learning bug and their reproduction.
By analyzing the user study results and the developers’ qualitative feedback, we gain new and valuable insights into the reproducibility of deep learning bugs. The user study results and feedback also provide information about the effectiveness of our recommended actions for bug reproduction. This enables us to deliver novel and actionable insights regarding the current state of deep learning bug reproducibility based on empirical evidence (Step 7, Fig. 1).
4 Study Findings
In this section, we present the findings of our study by answering three research questions as follows.
4.1 RQ1: Which edit actions are crucial for reproducing deep learning bugs?
We reproduced 85 bugs of different types and architectures to answer the first research question. Table. 5 briefly summarizes our reproduced bugs. Through our comprehensive reproduction process, we identify ten edit actions that are crucial for reproducing deep learning bugs. Table. 3 shows the identified actions from our qualitative analysis. We explain these actions as follows.
Input Data Generation (A) is one of the key edit actions for reproducing deep learning bugs. This action involves generating input data that simulates the data used for training. We generate the input data using the data shape and distribution provided by the reporter in a bug report. Our manual bug reproduction shows that 73% of our collected posts from Stack Overflow have the relevant data characteristics. For example, the following text from the Stack Overflow post (Issue #61781193) shows how the reporter might submit the input data distribution.
R: My training data has input as a sequence of 80 numbers in which each represent a word and target value is just a number between 1 and 3.
Using this information, we generated the random input data as follows and were able to reproduce the corresponding bug.
Neural Network Construction (A) was one of the most used edit actions during our bug reproduction. In this edit action, we construct a neural network based on the architecture provided by the reporter. Similar to the data characteristics, the information about the neural network is present in 65% of the reproducible issue reports. Using the neural network description from the issue reports, we were able to construct the models. For example, the following text from the Stack Overflow post (Issue #63204176) mentions the neural network but does not provide a code snippet for the network.
R: I am trying to write a simple multinomial logistic regression using mnist data.
Using this partial information, we constructed multiple multinomial logistic regression models, and despite the lack of relevant code examples, we successfully reproduced the corresponding bug.
Hyperparameter Initialization (A) is one of the core edit actions for reproducing deep learning bugs. As a part of this edit action, we initialize various hyperparameters (e.g., number of epochs, batch size, optimizer) for training the neural networks. Sometimes, the reporter did not provide these configurations, and we had to initialize them using default parameters to reproduce the bug. This is further evidenced by the fact that only 53% of our collected posts from Stack Overflow have information about the hyperparameters. For example, the following code snippet from the Stack Overflow post (Issue #31880720) shows a case where the hyperparameters were not initialized.
To reproduce the bug, we initialized the batch size with commonly-used values and the number of epochs ranging from 1 to 10. Since we initialized these hyperparameters, we ran the edited code snippet five times to confirm the effectiveness of our edit operation. We were able to reproduce the bug in all five iterations. We performed the steps mentioned above for all 37 bugs involving hyperparameter initialization.
Import Addition and Dependency Resolution (A) are one of the most common edit actions for all types of bugs. In this edit action, we analyze the code manually and determine the dependencies required for the code snippet to run. Then, we install the dependencies and manually import them to complete the code snippet. This edit action was also used to reproduce traditional software bugs, as reported by Mondal et al. msr2019mondal . In the Stack Overflow posts we collected, 47% of the issue reports lacked the required dependency information. This significant absence of the dependency information further highlights the need for our suggested edit action. For instance, the code snippet reported in the Stack Overflow post (Issue #50306988) was incomplete, and the dependency details were missing. Hence, we resolved the dependencies (keras, numpy and random) and added the required import statements to the code snippet. With the resolved dependencies, we were able to reproduce the bug successfully.
Logging (A plays a crucial role in the reproduction of different types of bugs. This action involves adding log statements to the provided code snippet to verify the reproduction of a bug. Let us consider the issue reported in the Stack Overflow question (Issue #70546468). The reporter provides a numpy array with shape (1, 3). The reporter then vertically concatenates (stack depthwise) multiple copies of this array to create tensor groups with shapes (2, 3) and (1, 3). After this vertical concatenation process, the reporter expects the tensor’s shape to be (2, None, 3). Unfortunately, the tensor’s shape was (2, None, None), according to the reporter. To verify the claim made by the reporter, we introduce a log statement within the code snippet. When we run the modified code snippet, we observe the shape of the stacked tensor to be (2, None, None), thereby confirming the reproduction and validity of the bug.
Obsolete Parameter Removal (A) is a vital edit action to enhance the reproducibility of bugs where the differences in the library and framework versions may lead to problems. In this edit action, we remove all the obsolete parameters not supported by the newer version of the library or framework. For example, the code snippet in the Stack Overflow post (Issue #65992364) has the following line of code.
However, according to the API documentation of PyTorch 1.6.0, the backend parameter was no longer supported. We removed the obsolete parameter and thus were able to reproduce the bug.
Compiler Error Resolution (A) is one of the extensively employed edit actions in reproducing bugs. In this edit action, we resolve the compiler errors to reproduce the bug. For example, in the Stack Overflow post (Issue #71514447), the reporter reports that the training loss is significantly increasing after every epoch. They suspect that the bug might be due to the incorrect computation of loss values. Furthermore, to help the developers reproduce the bug, they provide a detailed description of the bug and a complementary code snippet. However, the code snippet provided by the user does not compile, as the criterion function has not been defined. Hence, we replaced the function parameter with the default value of cross-entropy loss, as shown below.
Using the edit action above, we resolved the compiler errors and thus were able to reproduce the corresponding bug.
Dataset Procurement (A) is a critical edit action to reproduce bugs that require specific datasets. In this edit action, we analyze the issue report and attempt to procure the dataset mentioned in the report. For instance, in the Stack Overflow post (Issue #73966797), the reporter mentioned that they used the CIFAR-10, a well-known dataset for object detection. To reproduce the bug, we downloaded the dataset from its original source666https://www.cs.toronto.edu/ kriz/cifar.html, and using the dataset and the training code, we could reproduce the bug.
Downloading Models & Tokenizers (A) is one of the edit actions that was frequently used to reproduce bugs in Large Language Models and Transformer-based architectures. In this edit action, we download the pre-trained models and tokenizers for reproducing a bug. For example, in the Stack Overflow post (Issue #69660201), the reporter has provided the following information.
R: In the preprocessing layer, I’m using a BERT preprocessor from TF-Hub.
Based on the above information, we added the relevant URLs and configured the code to download the preprocessor and encoder. After downloading the preprocessor and encoder, we successfully reproduced the corresponding bug.
Version Migration (A) is one of the commonly used edit actions for all types of bugs. In this edit action, we replace the outdated or deprecated versions of a library or framework with the latest version to reproduce the bug. Mondal et al. msr2019mondal introduced a similar edit operation known as ‘Code Migration’ in their work. This operation often confirms the validity of the bug by reproducing it with the latest version. For example, the post on Stack Overflow (Issue #45711636) describes an issue with the CNN architecture constructed by the user. However, the code snippet provided in the issue uses Tensorflow 1.3.0. To reproduce this issue in TensorFlow 2.14.0, we carefully migrated the model built in TensorFlow 1.3.0 to the syntax of Tensorflow version 2.14.0. Specifically, we updated the Sequential, Conv2D and MaxPooling2D constructors to match the TensorFlow 2.14.0 API syntax. We also changed padding and pool size parameter schemes and removed the obsolete input shapes. Finally, we upgraded the model compilation and imported the required Tensorflow 2.14.0 modules. After modifying the code snippet, we could reproduce the bug successfully in our runtime environment.
4.2 RQ2: What information categories and edit actions are useful for reproducing specific types of deep learning bugs?
4.2.1 Identifying useful Information in Bug Reports
While reproducing deep learning bugs, we kept track of information from bug reports that helped us reproduce them. After reproducing 85 bugs successfully, we have identified five useful pieces of information that can improve the chance of reproducing a bug. We discuss these factors in detail below.
Data (F): Data is one of the essential factors in ensuring the reproducibility of deep learning bugs. Deep learning systems heavily rely on the data breck2019data , and reproducing deep learning bugs becomes easier with access to the original data. Data helps us reproduce the deep learning bugs by providing the exact sample inputs that trigger the erroneous behaviour. By understanding the training data distributions and ranges, we can reconstruct the original training environment, which is crucial for reproducibility. However, the issue reports often lack direct information about the data. To address this problem, we collect information about the data by extracting the shape of the data, data distribution, type of variables and their corresponding ranges from the issue description. Leveraging this information and our proposed edit actions, namely Input Data Generation (A) and Dataset Procurement (A), we can generate or obtain the necessary data to reproduce deep learning bugs. According to our investigation, 73% of our reproduced bugs contained information about the data (see Table. 6) in their issue reports. For example, in the Stack Overflow post (Issue #43464835), the reporter has provided the dimensions and sample row of the training dataset, as shown below.
R: I have a train dataset of the following shape: (300, 5, 720)
Sample Input: [[[ 6. 11. 389. …, 0. 0. 0.]]]]
Using this information and our proposed edit action Input Data Generation (A), we first generated a data frame of size (300, 5, 720) that contains values in the range from 0 to 400. Then, by leveraging the generated data and other useful information, we successfully reproduced the corresponding bug.
Model (F): The model architecture describes the components of a deep learning model and how they transform inputs into outputs. Understanding the model architecture is crucial for reproducing deep learning bugs, as it provides insights into the model’s components, connectivity, and the required neural network architecture. According to our investigation, 65% of our reproduced bugs contained information about the model architecture (see Table. 6) in their issue reports. However, the complete source code implementing the model architecture might not always be available in the issue reports. To overcome this limitation, we carefully gather information about a model’s architecture, i.e., the number of layers, layer properties and activation function from the issue description. Leveraging this information and our proposed edit action Neural Network Construction (A), we reconstruct the model and reproduce the deep learning bugs. For example, let us consider the following text from a Stack Overflow post (Issue #63204176) that mentions the use of multinomial logistic regression on the MNIST dataset, as shown below. Unfortunately, the reporter fails to provide the code snippet for the model.
R: I am trying to write a simple multinomial logistic regression using mnist data.
Despite the absence of the code snippet, the issue description provides useful hints about the model architecture and dataset. Using them and our edit action – Neural Network Construction (A), we successfully reproduced the bug.
Hyperparameters (F): Hyperparameters play a crucial role in controlling the learning process and model behaviour during training. They encompass parameters such as learning rate, batch size, number of epochs, optimizer, regularization techniques, and loss functions. The specific values chosen for these hyperparameters significantly impact a model’s performance, training process, and bug manifestation. Thus, reporting the complete set of hyperparameters is essential to help reproduce deep learning bugs. According to our investigation, 52% of our reproduced bugs contain information about the hyperparameters used in their issue reports. For the 48% of bug reports that did not include hyperparameter information, we used the Hyperparameter Initialization edit action (A) to reproduce the bugs by initializing the hyperparameters with default values. In the example Stack Overflow post (Issue #65993928), we can see how hyperparameters can play a crucial role in reproducing the bug.
Since the reporter was using a custom loss function, it was vital for them to share the value of constants and the formula used to calculate the loss. The reporter’s custom loss function, with constants 2 and 0.5 and its formula, was crucial for reproducing a bug in model training due to incorrect loss calculation, emphasizing the need to provide a complete set of hyperparameters.
Code Snippet (F): Code snippets are critical for reproducing bugs, as highlighted by the fact that 98% of professional developers consider them an essential component of bug reproducibility soltani2020significance . They include the data preprocessing, data splitting technique, code used for training the model, and implementation of the evaluation metrics. However, from our manual bug reproduction, we observe that even though code snippets are present in 79% of the bug reports, only 9.41% of them can be used verbatim for bug reproduction. To address this limitation, we use several of our proposed edit actions, such as Import Addition and Dependency Resolution (A), Logging (A), Obsolete Parameter Removal (A), Compiler Error Resolution (A) and Version Migration (A). These edit actions help us fix the errors in the code and make the code compileable and runnable.
To demonstrate the importance of submitting a complete code snippet, let us consider the Stack Overflow post with Issue #76186890. In this issue, a high-quality code snippet helps us reproduce the bug related to a complicated architecture (T5) without significant changes. The code snippet uses the T5 model from HuggingFace Transformers for text-to-text translation. It preprocesses the input text data, configures the target IDs to predict a specific answer, calculates loss and perplexity metrics, and trains the model. Having a complete and runnable snippet was helpful in reproducing the bug. It provides sufficient details like data preparation, loss calculation, model training, and evaluation to improve the ease of reproducing deep learning bugs, which were essential for bug reproduction.
Logs (F): Logs provide a real-time record of the model’s behaviour during training and inference. Traditional software logs consist of information such as event logs and stack traces, whereas deep learning systems consist of compiler error logs, training error logs, and evaluation logs Chen_Jiang_2022 . Sharing these logs is crucial for reproducing deep learning bugs as they allow us to verify if we can reproduce the same erroneous behaviour reported in the original issue. From our manual bug reproduction, we discover that 86% of our reproduced bugs contain the necessary logs in their issue reports. Such a high presence of logs in our dataset of reproduced bugs highlights the importance of logging in deep learning bug reproduction. By matching the logs from the original issue and the logs from reproduction on our local machines, we were able to confirm the presence of several bugs in the deep learning systems.
The Stack Overflow post (Issue #34311586) shows how logs can be used to confirm the presence of deep learning bugs. In this particular issue, the reporter shared the training logs and the code snippet. We modified the code snippet with our proposed edit actions to make it compilable and runnable. When executed, we observed the same anomalous behaviour in the training logs as described by the original reporter. This demonstrates the importance of sharing training and evaluation logs in the issue report. We also found them to be one of the most useful pieces of information to reproduce deep learning bugs.
| Factor | Data | Model | Hyperparameters | Code Snippet | Logs |
| Prevalence | 72.9% | 64.7% | 52.7% | 78.8% | 85.8% |
4.2.2 Relationship between Useful Information and Type of Bugs
During our manual analysis (Section 4.2.1), we identify useful information for reproducing deep learning bugs. We used the Apriori algorithm to determine the relationship between the type of bug and the information required to reproduce the bug. The insights from this analysis could serve two purposes – detecting missing information in submitted bug reports and formulating follow-up questions to obtain any missing information needed for reproducibility. Table. 7 summarizes our findings, and we discuss them in detail below.
Data: The accurate reproduction of bugs in deep learning systems heavily relies on the presence of data and its characteristics. They play a crucial role in reproducing two key categories of bugs - training bugs and tensor bugs.
For training bugs, the data has a confidence value of 80.77% according to our generated rules (check Table. 7). This high confidence highlights the significance of data in reproducing the numerous training issues that can manifest during model development. The details about the data, such as the number of samples, class distribution, feature distributions, data splitting ratios, and preprocessing steps, are instrumental in reproducing training bugs. For example, training a classification model on a highly skewed dataset is prone to overfitting. With the knowledge of the data distribution, we can recreate a representative dataset, which can help us reproduce the bugs, even if the actual dataset is not available.
For tensor bugs, data characteristics have an even higher confidence of 95.24%, according to our generated rules (Table. 7). This suggests that tensor attributes like shape, data type, sparsity, value ranges and origin are pivotal for reliably reproducing many bugs stemming from invalid dimensions or precision issues. For instance, bugs arising from tensor shape mismatches can emerge if the shape of the input data does not match the expected input shape. Furthermore, real-world data is often more vulnerable to human error and biases compared to synthetic data talwar2020evaluating . Therefore, comprehensive documentation and availability of these salient data characteristics are imperative for the reliable reproduction of the numerous training and tensor bugs in deep learning systems.
Model: A neural network model’s architecture and implementation details are crucial to reproducing model and API bugs. The model architecture has 70% and 72.71% confidence values for model and API bugs, respectively, according to our generated rules. These high values signify that access to model details is vital for reliably reproducing issues stemming from model capacity, connectivity, and API usage.
An access to the model architecture (e.g., layer types, layer connectivity, weight initialization schemes, etc.) can help us systematically reproduce model bugs. Insufficient learning capacity in specific layers and improper weight initialization might lead to exploding/vanishing gradients grosse2017lecture , whereas incorrect layer connectivity might lead to representational bottlenecks Tishby_Zaslavsky_2015 . These issues usually manifest as model bugs during training and inference. Thus, information about the model architecture can help us localize and reproduce such bugs.
For reproducing the API bugs, model architecture can provide fundamental context. Details like layer dimensions, bottlenecks, parallelization needs, and memory requirements show how the model interacts with the API yang2022comprehensive . Bottlenecks within a model, areas where data processing slows down, and the need for parallel processing to handle large-scale computations shape how APIs are utilized. Additionally, the varying memory requirements of different architectures impact how the model leverages the system’s resources via the API. This includes memory allocation for storing weights and activations and managing the flow of data through the network during operations like forward and backward propagation. Understanding the model architecture is vital for reproducing bugs, as it indicates whether issues stem from the model’s design or from its interaction with the API. For example, a bug might arise due to the model’s inability to handle certain operations efficiently, or it could be a result of the API not properly supporting specific architectural features. Hence, model architecture helps us reproduce the bugs triggered by incorrect usage of API. Therefore, the presence of model architecture allows the reproduction of both model bugs and API bugs in deep learning systems.
Code Snippet: Code snippets are invaluable for reproducing deep learning bugs since they isolate and encapsulate the core logic that triggers them. Developers can demonstrate and share the essence of buggy behaviour by creating a minimal reproducible example. Code snippets have very high confidence values of 92.3% for training bugs, 66.67% for tensor bugs, 70.0% for model bugs, and 81.8% for API bugs according to our generated rules. These high values across all bug types signify that code snippets are critical for reliably reproducing deep learning bugs.
Each code snippet may contain the relevant model, data processing, training and evaluation scripts. Developers can use such a snippet to recreate the bug-inducing steps. For example, a snippet may compactly capture just a few lines, mishandling tensor shapes or performing incorrect gradient calculations, eliminating any confounding factors and enabling developers to reproduce these bugs. Developers can systematically execute, analyze, and debug the code to reproduce various bugs, including training, tensor, model, and API bugs.
Logs: Logs play a fundamental role in deep learning by facilitating the reproduction and resolution of bugs. They comprehensively record every detail during model training, capturing information about various conditions, configurations, and events before the failures. Logs have high confidence values of 84.77% for training bugs, 85.71% for tensor bugs, 80% for model bugs, and 90.91% for API bugs according to our generated rules. Such high confidence levels across all bug types highlight that comprehensive log recording is important for the reliable reproduction of deep-learning bugs.
These logs include essential components such as hyperparameters, dataset characteristics, hardware specifications, framework versions, and random number generator seed values. The true power of logs lies in their ability to recreate past training runs precisely. Developers can isolate and reproduce the environment that led to the original bugs using the logged hyperparameters, such as batch size, learning rate schedules, and gradient clipping thresholds. Logs can also help developers reproduce specific deep-learning bugs. For example, logged random seeds can help the developers recreate a specific weight initialization or data batching order, which can then be used to reproduce training bugs. Similarly, logs can be used to reproduce Model, Tensor and API Bugs. Thus, logs are invaluable for accurately recreating conditions that result in errors and enabling the reproduction of deep learning bugs.
| Training | Model |
| Code Snippet (0.9231) | Logs (0.80) |
| Logs (0.8477) | Model (0.70) |
| Data (0.8077) | Code Snippet (0.70) |
| Tensor | API |
| Data (0.9524) | Logs (0.9091) |
| Logs (0.8571) | Code Snippet (0.8182) |
| Code Snippet (0.6667) | Model (0.7271) |
| Training | Model |
| Input Data Generation (0.6154) | Neural Network Definition (0.75) |
| Import Addition (0.4231) | Import Addition (0.55) |
| Version Migration (0.3477) | Hyperparameter Initialization (0.50) |
| Hyperparameter Initialization (0.3077) | Logging (0.45) |
| Compiler Error Resolution (0.2692) | Dataset Procurement (0.40) |
| Tensor | API |
| Import Addition (0.4545) | Input Data Generation (0.5714) |
| Input Data Generation (0.3636) | Hyperparameter Initialization (0.5238) |
| Hyperparameter Initialization (0.3636) | Import Addition (0.4762) |
| Dataset Procurement (0.2727) | Logging (0.4286) |
| Logging (0.1818) | Obsolete Parameter Removal (0.3333) |
4.2.3 Relationship between Edit Actions and Type of Bugs
After determining the association between the useful information needed and the type of bug, we derived the relationship between the type of bug and the edit actions. We use the Apriori algorithm to determine the relationship, as done earlier. Table. 8 summarizes our findings, and we discuss them in detail below.
Training Bug: Input data generation (61.54% confidence) is the most frequently used edit action for reproducing training bugs. However, our analysis (Table. 6) shows that 73% of these reports contain information about the data. Leveraging this information and our proposed edit action – Input Data Generation (A), we can reproduce the training bugs in deep learning systems. Furthermore, import addition (42.3% confidence) and version migration (34.7% confidence) are other edit actions which are often used to reproduce training bugs. Although 79% of bug reports contain code snippets, they are often incomplete. To reproduce the bug effectively, we thus need to complete these snippets by adding imports and migrating them to the latest versions of libraries and frameworks. Finally, hyperparameter initialization (30.77% confidence) and compiler error resolution (26.92% confidence) may also help reproduce training bugs by initializing the missing hyperparameters and resolving compile-time errors from the code snippet.
Model Bug: Model bugs are often reproduced by modifying a neural network’s architecture, as shown by the high confidence value (75%) for the edit action – neural network construction. Since model bugs are primarily caused by errors in the neural network architecture taxonomyRealFaults , reconstructing the neural network might be the first step to reproduce them. Additionally, import addition (55% confidence) and initializing hyperparameters (50% confidence) are typical edit actions to reproduce the model bugs. Additionally, logging can also help the reproduction of model bugs by recording important details like layer outputs, gradients, and parameter values during training. These recorded values can then be compared between buggy and bug-free models, which can confirm the presence of a bug in the model. Thus, model bugs can be reliably reproduced through several edits that construct the neural network architecture, initialize hyperparameters, add logging, and procure datasets.
Tensor Bug: Data is the most important information for reproducing tensor bugs, as shown by the strong correlation between Tensor Bug & Data (see Table. 7). This phenomenon can be attributed to the fact that tensor bugs primarily relate to the data taxonomyRealFaults . Since tensor bugs are so data-dependent, dataset procurement emerges as an important edit action to reproduce them. Bug reports reference the exact dataset used (if well-known) or describe the data type, shape and distribution. With this information, input data generation and dataset procurement can generate or procure the data required for the reproduction of the tensor bugs. Like other deep learning bugs, import addition (45.45% confidence) and hyperparameter initialization (36.36% confidence) also play a key role in completing the code snippet and reproducing the tensor bugs. Thus, tensor bugs can be reliably reproduced through several edits that procure the datasets, generate input data, add logging, import the required dependencies, and initialize the hyperparameters.
API Bug: Since API Bugs stem from an improper usage of application programming interfaces (APIs), they do not focus as heavily on the model training process. As a result, bug reports might lack detailed information about the input data or hyperparameters used. This encourages the use of common edit actions such as input data generation (57.14% confidence) and hyperparameter initialization (52.38% confidence) when reproducing API bugs. Additionally, code snippets in bug reports are often incomplete, necessitating edit actions like import addition (47.62% confidence) and logging (42.86% confidence) to trace the intermediate states and outputs of the API. Thus, API bugs can be reliably reproduced through several edit actions that generate input data, import the required dependencies, add logging, initialize hyperparameters and remove obsolete parameters.
4.3 RQ3: Do the suggested edit actions and information categories contribute to enhancing the reproducibility of deep learning bugs in practice?
We conduct a developer study to determine if our suggested edit actions and information help one reproduce deep learning bugs. We prepared four sets of bugs, each consisting of two different types of bugs (see Section 3.6). We randomly assigned one of the four bug sets to each participant and instructed them to reproduce the bugs in their assigned set. We divided the participants into the control and experimental groups using stratified random sampling (see Section 3.6). The developers in both the control and experimental groups reproduced the same bugs from our prepared set of bugs. The only difference was that the experimental group received hints based on our findings to help reproduce the bugs, while the control group did not receive these hints. Since both groups received randomly selected sets from the same pool of four identical bug sets, each bug was reproduced independently by at least five developers across both control and experimental groups.
Table. 9 shows the bug reproducibility rate of the control and experimental groups across different sets of bugs. We found that developers in the control group could reproduce 72.91% bugs without hints. In contrast, the developers in the experimental group could reproduce 95.83% of bugs with our hints - a 22.92% increase. This significant increase in reproducibility rate demonstrates that our identified edit actions and information improved developers’ ability to reproduce deep learning bugs.
| % of bugs successfully reproduced by Control group | % of bugs successfully reproduced by Experimental group |
% Increase |
| Set 1 |
75.00 |
100.00 |
25.00 |
| Set 2 |
66.66 |
83.33 |
16.66 |
| Set 3 |
83.33 |
100.00 |
16.66 |
| Set 4 |
66.66 |
100.00 |
33.33 |
| Average | 72.91 | 95.83 | 22.92 |
We also collect the qualitative feedback from the participants to better understand their pain points in bug reproduction. Their feedback provided new insights not covered in our original study. Notably, 31.78% of developers highlighted the lack of standardized debugging tools as the primary challenge in bug reproduction. Furthermore, the developers mentioned missing information (data, logs, hyperparameters, and dependencies), and lack of unit testing or version control in DL systems as the most challenging aspects of reproducing deep learning bugs, as shown below.
We also analyze how our suggested actions and information help the participants in bug reproduction. According to the qualitative responses, 40.91% of developers found our suggested edit actions to be helpful for reproducing their assigned bugs. 54.55% of developers report that our suggested hints about the useful information helped them narrow down where to look in the code. Thus, the qualitative feedback below highlights the benefits of our findings in real-life settings.
We also calculate the time the control and experimental groups took to reproduce their assigned bugs. This helps us assess our findings’ benefits in reducing the time required for deep learning bug reproduction. Table. 10 shows the time taken to reproduce bugs by the control and experimental groups. The experimental group outpaced the control group in bug reproduction across all sets, with the most notable difference in Set 4, where the experimental group was 30.16% faster. On average, the control group lagged behind the experimental group by 24.35% across all sets. These results demonstrate that our recommended edit actions and information categories enabled the experimental group to reproduce deep learning bugs much faster than the control group that did not receive this information.
| Average time taken by Control Group (seconds) | Average time taken by Experimental Group (seconds) | % Decrease | |
| Set 1 |
1437 |
1088 |
24.28 |
| Set 2 |
1803 |
1563 |
13.31 |
| Set 3 |
2548 |
1792 |
29.67 |
| Set 4 |
2165 |
1512 |
30.16 |
In short, the developer study demonstrates the real-world benefits of our findings and provides evidence that our findings can enhance the reproducibility of deep learning bugs.
5 Discussions
In this section, we first provide actionable insights about the reproducibility of deep learning bugs and recommendations for further research (see Section 5.1). We then describe a proof-of-concept tool, dlRepro, that we developed to recommend edit actions and useful information for reproducing deep learning bugs (see Section 5.2).
5.1 Reproducibility of Deep Learning Bugs
From our manual reproduction and user study, we observe that API, Model, and Tensor bugs are relatively more straightforward to reproduce. This behaviour can be explained by the fact that these bugs are more specific to the location in the code where they originate. For example, faulty input data usually triggers tensor bugs, whereas incorrect usage of a framework’s API causes API bugs. In contrast, training bugs cover multiple issues related to deep learning model training, and GPU bugs relate to GPU devices for deep learning and manifest across GPU interactions.
This behaviour is further supported by the reproducibility rates and efforts involved in reproducing different types of bugs. The reproducibility rates of Training bugs and GPU bugs were 89.65% and 42.85%, respectively. On the other hand, the reproducibility rates for API, Model and Tensor bugs were 81.81%, 88.46%, and 80.75%, respectively. Even though the training bugs could be reproduced reliably, the efforts involved in reproducing the training bugs were significantly more than those of other bugs. The average time to reproduce the API, Model and Tensor bugs was 45.5, 43.9 and 43.8 minutes, respectively. On the other hand, the average time to reproduce the Training bugs was 52.5 and 57.33 minutes, respectively. These statistics highlight that we need to put more effort into reproducing the Training and GPU bugs, which calls for further research into the specific nature of Training and GPU bugs.
To assist future research in the reproducibility of deep learning bugs, we provide directions for future research below:
Understanding Training Bug Reproducibility: Training bugs are the most common type of bug in deep learning systems, accounting for 52.5% of all bugs taxonomyRealFaults . Training bugs have high reproducibility rates but also require significant effort to reproduce. This behaviour presents an opportunity to understand better what factors make training bugs more reproducible or harder to reproduce. By analyzing the training procedures, model architectures, optimizers, hyperparameters, and other elements that either aid or hinder reproducibility, we can uncover insights to guide the diagnosis and repair of training bugs. There is also room to develop improved tools and methodologies explicitly focused on efficiently reproducing the nuanced nature of training bugs.
Analyzing the GPU Bug Reproducibility Gap: The lower reproducibility rate for GPU bugs highlights a gap in understanding the interactions between deep learning code and the underlying GPU hardware or drivers. By further studying irreproducible GPU bugs and quantifying the aspects that impede reproducibility, such as hardware differences, software dependencies, and environmental factors, we can work towards solutions to increase reproducibility. Opportunities exist to build infrastructure, leverage containerization or virtualization, and create testing tools to better control for sources of non-determinism that impact GPU bug reproduction.
Reproducible Testbeds for Deep Learning: Ultimately, the variability in reproducing different categories of deep learning bugs motivates the need for reproducible testbeds. Shared sets of reproducible deep learning bugs with associated test cases, model architectures, training configurations, dependencies, and environmental contexts will accelerate future research. Constructing such testbeds requires systematic characterization of how these factors influence reproducibility. Reproducible testbeds also support developing specialized techniques for efficiently reproducing and debugging deep learning bugs.
5.2 dlRepro: Proof-of-Concept Tool for Improving Bug Reproduction
From our user study (see RQ3), we observed that the lack of automated debugging tools is one of the most significant challenges in deep learning bug reproducibility. To aid the developers in reproducing deep-learning bugs, we created a proof-of-concept tool, dlRepro 777https://dlrepro.streamlit.app/ (Fig. 2). This tool takes in either a Stack Overflow question ID or a bug report as input. A bug report should include a bug description, code snippets demonstrating the issue, and relevant tags categorizing the bug. Based on the input provided, it recommends the useful information and edit actions needed to reproduce the deep learning bug. We discuss the methodology of the tool below.
Preparing the Bug Reports: Bug reports can be passed to the tool in two ways, which are described below:
Manual Entry: The user can provide the bug description, code snippet, and tags directly to the tool.
Stack Overflow: The user can input the Stack Overflow Question ID into the tool. Using this question ID, our tool utilizes the Stack Overflow API to retrieve the bug description, code snippet, and tags associated with that question.
Tag Analysis: We match the tags provided in the Stack Overflow question to our taxonomy of bug-related tags (Table 1). If there is a perfect match, we determine the bug type directly. However, if there is no match, we use a SentenceTransformer model called all-MiniLM-L6-v2 reimers2019sentence to find the semantic similarity between the tags in the post and our taxonomy tags for different bug types. This allows us to determine the most likely bug type based on the semantic similarity of the tags, even when there is no direct match.
Use case of our POC tool: To demonstrate how our tool can be used to reproduce deep learning bugs, let us consider the Stack Overflow Issue #54028667. This issue 888https://stackoverflow.com/q/54028667 is caused by an incorrect usage of the scikit-learn API. The tags for this issue are ‘python’, ‘scikit-learn’, ‘gridsearch’, ‘gridsearch-cv’. Since none of these tags exactly match the tags from our analysis (Table. 1), the tool computes the semantic similarity to determine the type of bug. Based on semantic similarity values, the tool determines the type of bug to be API. After determining the type of bug, the tool consults our findings from RQ2 (Table. 7 and Table. 8). It presents useful information and key edit actions which can be used to reproduce the bug. The outputs from the tool are shown in Fig. 3 and Fig. 4.
The bug report contained all the information required for reproduction. Hence, by using the recommended edit actions and the complementary code snippets, we could reproduce the bug in 23 minutes, which shows the effectiveness of our findings and the tool. As the bug contained the required information, other people could also reproduce the issue, and the question received an accepted answer within 12 hours. Although our dataset consisted of posts containing ‘tensorflow’, ‘pytorch’, and ‘keras’, we could reproduce a bug related to ‘scikit-learn’. This highlights that our findings and tool are applicable beyond the specific libraries used to build the original dataset. Reproducing the ‘scikit-learn’ bug shows that our approach works even when applied to a library not seen during initial development and testing, highlighting the generalizability of our findings and tool.
Extending and Enhancing Usability of dlRepro: While our study focused on exploring and improving the state of deep learning bug reproducibility, dlRepro was developed as a proof-of-concept tool to demonstrate how our findings can serve as a step towards automated reproduction of deep learning bugs. The tool’s flexible architecture enables continuous offline updates to the underlying algorithm for tag similarities, information category mappings, and edit recommendations without changes to the user-facing interface. It also encourages broader deep-learning developer community contributions. Through this collaborative access model, experts could voluntarily expand the tool’s capabilities by sharing new insights uncovered. Such community participation can lead to gradual, real-world refinements and expansion of both the dlRepro tool and the taxonomy or tags that empower our tool. We have also open-sourced dlRepro’s code to enable direct extensions by the community dlReproCitation , further showcasing the potential for collaborative improvement. While demonstrating the applicability of our analysis, dlRepro establishes a framework for ongoing communal efforts to enhance deep learning bug reproducibility through collective intelligence.
5.3 Comparison with ChatGPT
To assess the effectiveness of our POC tool in deep learning bug reproducibility, we conducted a comparison with ChatGPT, considering the widespread usage of large language models. To do this, we provided ChatGPT with a bug report, supplementary information, and a specific query, as shown in Fig. LABEL:fig:chatGPTQuery. To evaluate ChatGPT’s capability to reproduce deep learning bugs, we supplied it with the identical bug report, code snippets, and other details utilized in our POC tool’s use case (Section 5.2).
The query included an instruction asking ChatGPT for help reproducing the bug, while the query context included the detailed bug description and corresponding code snippet. By providing this additional context, we aimed to enhance ChatGPT’s understanding of the bug. We wanted the ChatGPT to reproduce the bug.
Upon requesting ChatGPT to reproduce the bug, the model provided a potential solution to the problem. Although this solution could potentially resolve the bug, it might not be enough to reproduce the bug, which was our original goal. Unlike ChatGPT’s response, our findings allowed us to reproduce the bug successfully by combining edit actions and useful information, as shown in Section 5.2.
This case study serves as a compelling demonstration of the practicality and value of our findings, even in the era of language models. It highlights how our approach can offer unique insights and solutions that are not readily apparent through language models alone. By leveraging our findings and language models’ capabilities, we can maximize the effectiveness of bug reproducibility techniques and drive advancements in this field.
6 Threats to Validity
Threats to external validity relate to the generalizability of our findings. We have reproduced multiple bugs from five different types to mitigate this threat. Moreover, we have reproduced bugs from 14 different architectures, ranging from Logistic Regression, CNN, Transformers and BERT-based architectures. Finally, we have also reproduced the bugs from the past six years (2017-2023), and our findings align with and extend the previous findings of software bug reproducibility msr2019mondal , possibly indicating the generalizability of our study.
Threats to internal validity pertain to experimental errors and confounding variables during the reproduction of bugs. The possibility of introducing new bugs into the code exists during bug reproduction. We have implemented several strategies and precautions to mitigate these threats in our experimental design. For instance, bug reproduction attempts are carried out in a controlled and isolated environment. We refrain from using hardcoded or fixed values in our code snippets to prevent potential errors during bug reproduction. For instance, during the bug reproduction, we utilize random hyperparameter initialization, construct multiple neural networks, and generate randomized input data that aligns with the original distribution. This methodology mitigates the influence of unforeseen variables on the outcomes of our experiments. Following bug reproduction, we undertake rigorous manual analysis to identify any discrepancies or anomalies in the outcomes. Another threat to validity is the incomplete or incorrect information in the bug report, which may lead to incorrect conclusions about the reproducibility of deep learning bugs. To mitigate this threat, we apply extensive filtration based on the criteria suggested by Moravati et al. defects4ml and Humbatova et al. taxonomyRealFaults , which leads to a clean and generalizable dataset.
Threats to construct validity relate to the use of appropriate metrics for evaluating the results of this user study. We used the reproducibility rate and time to reproduce the deep learning bugs as the metrics to evaluate the benefits of our findings. These metrics directly measure how easily the bugs can be reproduced, which aligns with our goal of making deep learning bugs more reproducible. To ensure these metrics were accurately measured, we had clear criteria for determining whether a bug was successfully reproduced. We also recorded precise timestamps for how long it took participants to reproduce bugs. This helped mitigate threats related to measurement errors or subjective assessments of the metrics. Another threat to the validity might be the accuracy of tags used to categorize the bugs. The incorrect categorization of bugs might lead to incorrect conclusions for different bugs. To mitigate this threat, we only select the tags that are present in the taxonomy and the sub-taxonomies defined by Humbatova et al. taxonomyRealFaults . For example, we specifically chose the “loss-function” as one of the tags to distinguish the training bugs since the Loss Function falls within the sub-taxonomy of Training bugs. Thus, the threats to construct validity were effectively mitigated.
7 Related Work
There have been several studies that focus on the aspects of reproducibility of software bugs msr2019mondal ; mondal2022reproducibility ; mondalreproducibility , or focus on why some bugs cannot be reproduced icsme2020rahman ; rahman2022works . Many studies attempt to learn the nature of deep learning bugs tfProgramBugs ; islamfse19 ; taxonomyRealFaults ; tarek2022icsme , and how they can be localized automatically nerdbug ; yan2021exposing ; zhang2020detecting . A few studies also attempt to learn about the state of reproducibility of deep learning in software engineering trainingReproducibleDL ; reproducibilityDLTSE , and reproduce deep learning bugs as a part of benchmark dataset creation defects4ml . Unfortunately, only a little research has been done to understand the challenges in reproducing deep learning bugs and how we can improve their reproducibility.
Mondal et al. msr2019mondal extensively investigated the reproducibility of programming issues reported on Stack Overflow. They identified several key edit actions required to reproduce programming errors and traditional software bugs. While their work is a source of inspiration, it does not provide the edit actions for reproducing deep learning bugs. Since their dataset is constructed in Java programming language with no questions related to deep learning, their findings might not apply to deep learning bugs.
Chen et al. trainingReproducibleDL proposed a unified framework to train reproducible deep learning models. They also provide reproducibility guidelines and mitigation strategies for conducting a reproducible training process for DL models. However, the primary focus of their study was on the reproducibility of deep learning models, not deep learning bugs. Furthermore, the guidelines for deep learning models might not always extend to deep learning bugs.
Moravati et al. defects4ml constructed the faultload benchmark containing deep learning bugs, defects4ML. As a part of their benchmark dataset, they reproduced 100 bugs as a part of the reproducibility criterion for benchmark datasets. While their study was the first to reproduce deep learning bugs actively, they did not report the techniques and actions used to reproduce bugs. Furthermore, they did not report the useful information from the issue reports, which helped them reproduce the bugs.
Unlike many earlier studies above, we conduct an extensive empirical study to understand the challenges of deep learning bug reproduction and how we can improve the reproducibility of deep learning bugs. We construct a dataset of 668 bugs and manually reproduce 85 bugs, spanning three frameworks, five types of bugs, and 14 architectures. We not only (1) define ten edit actions which can be used to reproduce deep learning bugs but also (2) explore the associations among the type of bugs, the information required to reproduce them, and the specific edit actions that can be used to reproduce them, which makes our work novel. Our findings are also generalizable due to the diversity of our dataset. To ensure transparency and reproducibility, we have made our dataset and replication package publicly available999https://github.com/mehilshah/Bug_Reproducibility_DL_Bugs.
8 Conclusion & Future Work
The reproducibility of deep learning bugs is a significant challenge for software practitioners, given that these bugs can lead to severe consequences. In this paper, we conduct an empirical study by reproducing 85 deep learning bugs and identifying ten key edit actions that could reproduce these bugs. Furthermore, we investigate the relationships among bug types, information categories, and edit actions. To validate our findings, we conducted a user study with 22 participants. The developers equipped with our recommended edit actions and information could reproduce 22.92% more bugs using 24.35% less time compared to the developers without such hints. This demonstrates the real-world value of our findings in improving the reproducibility of deep learning bugs. Finally, we provide actionable insights about the state of deep learning bugs that could improve the reproducibility of deep learning bugs. To demonstrate the practical use of our findings, we built a proof-of-concept tool, dlRepro, that takes bug reports and recommends the edit actions and useful information for deep learning bugs. Moreover, we also conduct a comparison with ChatGPT, which fails to reproduce the bug, while our methodology succeeds. This demonstrates the real-world value of our findings in improving the reproducibility of the deep learning bugs. For future work, we plan to expand our study to additional deep learning frameworks and model types. We will also improve dlRepro with more advanced natural language processing capabilities to extract key information from bug reports. Finally, we aim to integrate dlRepro into developer tools and workflows to provide automated reproducibility support during deep learning development. Furthermore, our findings open up the ability to create automated systems for reproducing deep learning bugs. By training machine learning models on our dataset of reproduction actions matched to bug reports, we can build AI agents that take bug reports as input and automatically generate reproducibility scripts, similar to the work by White et al. reproducibleBugReports . This would significantly reduce the manual effort required to reproduce bugs and enable large-scale debugging and testing of deep learning systems.
Data Availability Statement (DAS)
All the data generated or analyzed during this study are available in the GitHub repository to help reproduce our results replicationPackage .
Conflict of Interest
The authors declare that they have no conflict of interest.